If f(x) = log10 x, find f(1), f(10), f(5)


1
Expert's answer
2022-06-07T14:49:03-0400

f(1) = log₁₀ (1) = 0

f(10) = log₁₀ (10) = 1

f(5) = log₁₀ (5) = log₁₀ (10/2) = log₁₀ (10) - log₁₀ (2) = 1 - 0.301 = 0.699